# Can Large Language Models Invent Algorithms to Improve Themselves?

### v3

@@ Abstract
Large Language Models (LLMs) have shown remarkable performance improvements and are rapidly gaining adoption in industry. However, the methods for improving LLMs are still designed by humans, which restricts the invention of new model-improving algorithms to human expertise and imagination. To address this, we propose the Self-Developing framework, which enables LLMs to autonomously generate and learn model-improvement algorithms. In this framework, the seed model generates, applies, and learns model-improving algorithms, continuously improving both the seed model and the algorithms themselves. Among model-improving strategies, we focus on model merging algorithms. In mathematical reasoning tasks, Self-Developing discovers novel merging strategies and outperforms human-designed methods. On GSM8k, the discovered algorithms improve the seed model by 6% and surpass human-designed methods by 4.3%. Moreover, they exhibit strong transferability, achieving a 7.4% performance gain on out-of-domain models. These results suggest that LLMs can autonomously develop effective model-improvement techniques beyond human intuition.
Can Large Language Models Invent Algorithms to Improve Themselves?
Yoichi Ishibashi NEC yoichi-ishibashi@nec.com                        Taro Yano NEC taro_yano@nec.com                        Masafumi Oyamada NEC oyamada@nec.com
@@ Results
Self-improving
The concept of self-improving artificial intelligence was proposed by Minsky (1966) and Good (1965), and later formalized by Schmidhuber (2003). With the rapid development of LLMs, the community has shifted towards practical implementations of self-improvement Huang et al. (2023). Many recent self-improvement approaches primarily focus on the generation of fine-tuning with self-generated training data Yuan et al. (2024); Gülçehre et al. (2023); Zhang et al. (2024); Wang et al. (2023); Xu et al. (2024a). Their methods do not generate or learn the improvement strategies themselves. Additionally, agents that modify outputs based on feedback from the environment have been proposed Madaan et al. (2023); Shinn et al. (2023); Ishibashi and Nishimura (2024), but these are different from our goal of improving the LLM itself.
Algorithm Generation using LLMs
Code generation by LLMs Jiang et al. (2024) has been proposed for various applications, such as solving reasoning problems Chen et al. (2023) and generating agent actions Wang et al. (2024). Focusing on LLMs’ code generation capabilities, several approaches have been suggested where LLMs generate and execute code to enhance their own abilities. For example, Lu et al. (2024) propose a method where LLMs themselves are used to discover loss functions for preference optimization. Zelikman et al. (2023) propose a method to improve a code that makes structured calls to a LLM. These methods generate improvement algorithms using LLMs. Unlike their approaches, we not only generate model-improvement algorithms but also enhance the LLM that generates these algorithms. Another major difference is that we do not rely on external LLMs, like GPT-4 OpenAI (2023), other than the seed model, for algorithm generation. Our work builds upon these foundations but uniquely focuses on the autonomous improvement of both the algorithms and the model generating them, without relying on external models, thus pushing the boundaries of self-improving AI systems.

### v5

@@ Abstract
Large Language Models (LLMs) have achieved remarkable capabilities, yet their improvement methods remain fundamentally constrained by human design. We present Self-Developing, a framework that enables LLMs to autonomously discover, implement, and refine their own improvement algorithms. Our approach employs an iterative cycle where a seed model generates algorithmic candidates as executable code, evaluates their effectiveness, and uses Direct Preference Optimization to recursively improve increasingly sophisticated improvement strategies. We demonstrate this framework through model merging, a practical technique for combining specialized models. Self-Developing successfully discovered novel merging algorithms that outperform existing human-designed algorithms. On mathematical reasoning benchmarks, the autonomously discovered algorithms improve the seed model’s GSM8k performance by 6% and exceed human-designed approaches like Task Arithmetic by 4.3%. Remarkably, these algorithms exhibit strong generalization, achieving 7.4% gains on out-of-domain models without re-optimization. Our findings demonstrate that LLMs can transcend their training to invent genuinely novel optimization techniques. This capability represents a crucial step toward a new era where LLMs not only solve problems but autonomously develop the methodologies for their own advancement.
Can Large Language Models Invent Algorithms to Improve Themselves?:
Algorithm Discovery for Recursive Self-Improvement through Reinforcement Learning
Yoichi Ishibashi NEC Corporation yoichi-ishibashi@nec.com                        Taro Yano NEC Corporation taro_yano@nec.com                        Masafumi Oyamada NEC Corporation oyamada@nec.com
@@
Figure 1: The concept of our Self-Developing framework. The algorithm factory (LLM) generates model-improvement algorithms as executable code and evaluates their effectiveness (/). High-performing algorithms serve as positive examples and low-performing ones as negative examples to train the algorithm factory via reinforcement learning, enabling it to generate better algorithms in subsequent iterations. Simultaneously, these improved algorithms are applied to the seed model to create enhanced variants, thus improving both the algorithm generation capability and the target models.
@@ Results
Recursive self-improvement
The concept of self-improving artificial intelligence was proposed by Minsky (1966) and Good (1965), and later formalized by Schmidhuber (2003). With the rapid development of LLMs, the community has shifted towards practical implementations of self-improvement Huang et al. (2023). Many recent self-improvement approaches primarily focus on the generation of fine-tuning with self-generated training data Yuan et al. (2024); Gülçehre et al. (2023); Zhang et al. (2024); Wang et al. (2023); Xu et al. (2024a). Their methods do not generate or learn the improvement strategies themselves. Additionally, agents that modify outputs based on feedback from the environment have been proposed Madaan et al. (2023); Shinn et al. (2023); Ishibashi and Nishimura (2024), but these are different from our goal of improving the LLM itself. Regarding research on automating LLM development, Yano et al. (2025) have worked on the automatic construction of post-training pipelines. They employ LLM-based agents to automatically search for combinations of techniques such as SFT and model merging. In contrast, our focus is on the automatic discovery of model improvement algorithms.
Algorithm discovery using LLMs
Code generation by LLMs Jiang et al. (2024) has been proposed for various applications, such as solving reasoning problems Chen et al. (2023) and generating agent actions Wang et al. (2024). Code generation has also achieved significant results in scientific discovery. AlphaEvolve Novikov et al. (2025) has demonstrated practical achievements in mathematical discovery and system optimization by combining evolutionary computation with LLMs. In the context of recursive self-improvement, Lu et al. (2024) propose a method for discovering loss functions for preference optimization using LLMs themselves, and Zelikman et al. (2023) propose a method for improving code that makes structured calls to LLMs. However, existing methods have several important limitations: (1) No generator improvement: The algorithm generator itself is not improved and lacks the ability to learn from generated algorithms to produce better algorithms, (2) External dependency: They rely on more powerful external models such as GPT-4 OpenAI (2023). Self-Developing overcomes these limitations by continuously training the algorithm generator (Algorithm Factory) itself using DPO. By using preference data constructed from pairs of high-performing and low-performing algorithms, the algorithm generator learns the characteristics of effective algorithms and becomes capable of efficiently generating higher-quality algorithms over time. Furthermore, it achieves recursive self-improvement without depending on external models other than the seed model. We are the first to achieve improvement of both the algorithms and the LLM that generates them in the context of recursive self-improvement. To the best of our knowledge, we are the first to recursively improve an algorithm-generating LLM by training it to generate better algorithms.